Grade 11Physical Chemistry

the number of atoms in 4.25g of NH3 is approximately.

Himanshu
8 Years agoGrade 11

1 Answer

Jobin C J
8 Years ago
Answer is 6.022×10^23StepsNo.of Moles=Given Mass / Molar MassGiven Mass =4.25 g =17/4 gMolar Mass of NH3 = 17g So No.of Moles =17/4 ÷ 17= ¼Therefore No.of atoms=No.of Moles×6.022×10^23×AtomicityAtomicity of NH3 is 4i.e. No.of atoms=¼ × 6.022×10^23 × 4 = 6.022×10^23